OpenStudy (anonymous):

11 1/3 divided by 2 5/6

OpenStudy (anonymous):

make the mixed fractions into improper fractions: (34/3) / (17/6) then to divide fractions you make the denominator fraction to its reciprocal and multiply (34/3) x (6/17) = 204/51 = 4

OpenStudy (anonymous):

34/3 x 6/17 204/51 4